OpenStudy (anonymous):

What are the mean and the mode of the following set of data: 5, 12, 1, 5, 7

OpenStudy (anonymous):

mode is 5 since it repeated the most

OpenStudy (anonymous):

To find the mean, you add up all the numbers and then divide it by 5 because there are five numbers.

OpenStudy (anonymous):

mean is 5 + 12 + 1 + 5 + 7 / 5 = 6

OpenStudy (anonymous):

Mean is like the average so Add them all up and divide by the # of values there are 5+ 12 + 1 + 5 + 7 / 5 =30/5 = 6 So mean is 6 And mode is just the # that occurs the most in a data set